The width of a rectangle is 1 ft less than the length. the area is 2 ft^2. Find the length and the width.
Answer by checkley77(12844)

W=L-1
WL=2
(L-1)L=2
L^2-L=2
L^2-L-2=0
(L-2)(L+1)=0
L-2=0
L=2 ANS. FOR THE LENGTH.
W=2-1=1 ANS. FOR THE LENGTH.
PROOF:
2*1=2
2=2